OpenStudy (anonymous):

Please help me... Word Problem Solving A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

Let the distance b/n the two place be x and the velocity of the plane be v. On the way to Denver the equation is x/5 = v + 40 And the to Orland the equation will be x/6 = v - 40 Now solve for V you will get v = 40

OpenStudy (anonymous):

Thank you!